Please help! Place the following points in the a Number Line: (9/7), (-25/6), -13, (-25/8), 0, (25/6), (25/8), 13, (-9/7), (25/2)
Answer by Edwin McCravy(20062)

(9/7), (-25/6), -13, (-25/8), 0, (25/6), (25/8), 13, (-9/7), (25/2)

Change them all to approximately decimals, so you can know their sizes

9/7 = 1.29
-25/6 = -4.17
-13 
-25/8 = -3.125
0 
25/6 = 4.17
25/8 = 3.13
13 
-9/7 = -1.29
25/2 = 12.5

Arrange them in order from smallest to largest:

  -13 = -13
-25/6 =  -4.17
-25/8 =  -3.125
 -9/7 =  -1.29
    0 =   0
  9/7 =   1.29
 25/8 =   3.13
 25/6 =   4.17
 25/2 =  12.5
   13 =  13

Make a dot on the number line corresponding to each number and write the number
below or above that dot.  You can just estimate the fraction of a distance
between the integers as the decimal part indicates:



Edwin